OpenStudy (anonymous):

integrate: -0.5537e^-0.03767t

OpenStudy (tkhunny):

integrate: 7e^6t

OpenStudy (unklerhaukus):

\[\int ae^{bx}\,\mathrm dx=a\int e^{bx}\,\mathrm dx\\ \qquad\qquad=a\frac{e^{bx}}b+c\] \[\int-0.5537e^{-0.03767t}\,\mathrm dt\\=-0.5537\int e^{-0.03767t}\,\mathrm dt\\=\]